# Analytical description for light propagation with the source inside the black hole photon sphere

### v1

@@ I Introduction
In the spacetime of a black hole, the deflection of photons gives rise to the phenomenon of gravitational lensing, which serves as a powerful tool for studying the physical properties of the black hole and has been extensively studied in the literature Perlick (2004); Bozza (2010); Cunha and Herdeiro (2018). In the asymptotically flat region far away from the black hole, photons get weakly deflected and two images may be generated due to gravitational lensing Keeton and Petters (2005). These images have the potential to be confirmed by the monitoring of stars near Sgr A* in our galactic center by GRAVITY and then play a crucial role in understanding the structure of the spacetime Keeton and Petters (2006); GRAVITY Collaboration (2018). In the vicinity of the black hole event horizon, photons can be strongly deflected and gravitational lensing generates infinite relativistic images and photon rings that are important observational targets for the next-generation Event Horizon Telescope or future space-borne very long baseline interferometry Akiyama et al. (2019) (Event Horizon Telescope Collaboration); Akiyama et al. (2022) (Event Horizon Telescope Collaboration); Johnson et al. (2020); Vincent et al. (2022); Paugnat et al. (2022), enabling our in-depth exploration of the strong-field physics Gralla et al. (2019); Wang et al. (2019); Lu and Xie (2019); Gao and Xie (2021); Zhang and Xie (2022); Aratore and Bozza (2021); Gao and Xie (2024).
@@
To investigate the deflection of photons emitted by the source inside the photon sphere in the black hole spacetime, either numerical or analytical methods can be employed Perlick and Tsupko (2022). Numerical methods typically involve forward or backward ray tracing algorithms, which require direct integration of the geodesic equations of photons Zhou et al. (2024). Alternatively, analytical approaches usually obtain approximate solutions for key physical quantities like the deflection angle and travel time through series expansion Keeton and Petters (2005); Bozza (2002); Bozza and Mancini (2004); Bozza and Scarpetta (2007); Jia and Huang (2021); Liu and Jia (2021); Poutanen (2020); Claros and Gallo (2024), and can reveal the underlying physics behind the practical observations more effectively. Among all the analytical methods, the one based on the photons’ impact parameter can describe the strong deflection of photons near the photon sphere in a completely analytical manner Jia and Huang (2021), which has been successfully applied in the case where the source is outside the photon sphere. In this work, we will generalize this method to the case where the source is inside the photon sphere.
@@
If the photon is emitted by a source far away from the black hole, the corresponding deflection angle is determined by the closest distance of approach (i.e., the turning point in the photon’s trajectory). The closer the closest distance to the black hole is, the larger the deflection angle will be. When the closest distance attains a critical value, the deflection angle diverges Bozza and Scarpetta (2007). This critical value, denoted as $r_{\mathrm{m}}$, satisfies the following equation Virbhadra and Ellis (2000); Claudel et al. (2001)
@@
    Photons with $u>u_{\mathrm{m}}$ and $r_{\mathrm{S}}>r_{\mathrm{m}}$. As shown by the red dotted line, these photons are emitted by the source outside the photon sphere, and their initial direction can be either radially inward or outward. For the inward ones, they can be received by the observer after being weakly deflected when $u$ is much larger than $u_{\mathrm{m}}$ Keeton and Petters (2005); Huang and Jia (2020) or being strongly deflected when $u$ is slightly larger than $u_{\mathrm{m}}$ Bozza (2002); Jia and Huang (2021). For the outward ones, they can also be observed if they are directly facing the observer.
@@
    Photons with $u>u_{\mathrm{m}}$ and $r_{\mathrm{S}}<r_{\mathrm{m}}$. They originate from the region inside the photon sphere and outside the event horizon. Given that $u>u_{\mathrm{m}}$, these photons feature turning points in their trajectories and will eventually fall into the horizon Bozza and Scarpetta (2007), as depicted by the red solid line.
@@
For photons of type IV propagating in the black hole spacetime, the change in the azimuthal angle and the travel time can be respectively obtained as
@@ III.1 Comparison with previous works and numerical results
The leading-order analytical approximation of $\Delta\phi$ for photons emitted by the source outside the photon sphere is given by Bozza and Scarpetta (2007)
@@
where $a$ and $b$ are the strong deflection coefficients that are directly associated with the metric functions $A(r_{\mathrm{m}}),B(r_{\mathrm{m}})$ and $C(r_{\mathrm{m}})$.
Eq. (33) has been demonstrated to be applicable to photons emitted by sources at any distance outside the event horizon Bozza and Scarpetta (2007). In comparison, Eq. (27) is only applicable for the case where the source is inside the photon sphere. However, our approximation gives the higher-order corrections that are omitted by Ref. Bozza and Scarpetta (2007), and is thus more accurate.
@@ IV.1.2 Comparison with numerical results
From the top panel of Fig. 4 we can observe that when $N=5$, the value of the time delay $\Delta T$ is consistent with $2(\pi-\phi_{\mathrm{S}})u_{\mathrm{m}}$ predicted by Eq. (45). However, $\Delta T$ is not available for $\phi_{\mathrm{S}}=0$. This is because the impact parameter $u|_{n=0}$ and its corresponding $\epsilon$ defined by (13) are both equal to 0, which causes Eq. (37) with $N=5$ to diverge. Such a problem always exists, but $\Delta T$ can be evaluated in the case of $N=0$ due to the relation (41). When $N=5$, the relation (41) is corrected by higher-order terms, and one thus has to calculate $\Delta T$ from Eq. (37) directly.
@@
Furthermore, when calculating the time delay, we have made the assumption that the source, lens and observer lie on a fixed plane, which is just a simple scenario. In more realistic contexts, for instance, when the source maintains a plunging orbit inside the photon sphere, the plane on which the source, lens and observer are located will change over time. How the resulting time delay evolves along the plunging orbit will be our next move.
Finally, black hole mimickers such as ultracompact objects can have an internal structure Cardoso and Pani (2019), and photons emitted by the source inside their photon sphere may exhibit different time delay characteristics. How to distinguish black holes and their mimickers through the time delay also deserves future work.

### v2

@@ I Introduction
In the spacetime of a black hole, the deflection of photons can give rise to the phenomenon of gravitational lensing, which serves as a powerful tool for studying the physical properties of the black hole and has been extensively studied in the literature Perlick (2004); Bozza (2010); Cunha and Herdeiro (2018). In the asymptotically flat region far away from the black hole, photons get weakly deflected and two images may be generated due to gravitational lensing Keeton and Petters (2005). These images have the potential to be confirmed by the monitoring of stars near a supermassive black hole by GRAVITY and to play a crucial role in understanding the structure of the spacetime Keeton and Petters (2006); GRAVITY Collaboration (2018). In the vicinity of the black hole event horizon, photons can be strongly deflected and gravitational lensing generates infinite relativistic images and photon rings that are important observational targets for the next-generation Event Horizon Telescope or future space-borne very long baseline interferometry Akiyama et al. (2019) (Event Horizon Telescope Collaboration); Akiyama et al. (2022) (Event Horizon Telescope Collaboration); Johnson et al. (2020); Vincent et al. (2022); Paugnat et al. (2022), enabling our in-depth exploration of the strong-field physics Gralla et al. (2019); Wang et al. (2019); Lu and Xie (2019); Gao and Xie (2021); Zhang and Xie (2022); Aratore and Bozza (2021); Gao and Xie (2024).
@@
To investigate the deflection of photons emitted by the source inside the photon sphere in the black hole spacetime, either numerical or analytical methods can be employed Perlick and Tsupko (2022). Numerical methods typically involve forward or backward ray tracing algorithms, which require direct integration of the geodesic equations of photons Zhou et al. (2024). Alternatively, analytical approaches usually obtain approximate solutions for key physical quantities like the deflection angle and travel time through series expansion Keeton and Petters (2005); Bozza (2002); Bozza and Mancini (2004); Bozza and Scarpetta (2007); Jia and Huang (2021); Liu and Jia (2021); Poutanen (2020); Claros and Gallo (2024); Igata (2025), and can reveal the underlying physics behind the practical observations more effectively. Among all the analytical methods, the one based on the photons’ impact parameter can describe the strong deflection of photons near the photon sphere in a completely analytical manner Jia and Huang (2021), which has been successfully applied in the case where the source is outside the photon sphere. In this work, we will generalize this method to the case where the source is inside the photon sphere.
@@ II.1 Classification of different photons in the black hole spacetime
If the photon is emitted by a source far away from the black hole, the corresponding deflection angle is determined by the closest distance of approach (i.e., the turning point in the photon’s trajectory). The closer the closest distance to the black hole is, the larger the deflection angle will be. When the closest distance attains a critical value, the deflection angle diverges Claudel et al. (2001); Bozza and Scarpetta (2007). This critical value, denoted as $r_{\mathrm{m}}$, satisfies the following equation Virbhadra and Ellis (2000); Claudel et al. (2001); Adler and Virbhadra (2022)
@@
    Photons with $u>u_{\mathrm{m}}$ and $r_{\mathrm{S}}>r_{\mathrm{m}}$. As shown by the red dotted line, these photons are emitted by the source outside the photon sphere, and their initial direction can be either radially inward or outward. For the inward ones, they can be received by the observer after being weakly deflected when $u$ is much larger than $u_{\mathrm{m}}$ Keeton and Petters (2005); Huang and Jia (2020) or being strongly deflected when $u$ is slightly larger than $u_{\mathrm{m}}$ Claudel et al. (2001); Bozza (2002); Jia and Huang (2021). For the outward ones, they can also be observed if they are directly facing the observer.
@@
    Photons with $u>u_{\mathrm{m}}$ and $r_{\mathrm{S}}<r_{\mathrm{m}}$. They originate from the region inside the photon sphere and outside the event horizon. Given that $u>u_{\mathrm{m}}$, these photons feature turning points in their trajectories and will eventually fall into the horizon Claudel et al. (2001); Bozza and Scarpetta (2007), as depicted by the red solid line.
@@ II.2 Light propagation with the source inside the photon sphere
For photons of type IV propagating in the black hole spacetime, the change in the azimuthal angle and the travel time can be respectively obtained as Virbhadra and Keeton (2008)
@@ III.1 Comparison with previous works and numerical results
The leading-order analytical approximation of $\Delta\phi$ for photons emitted by the source inside the photon sphere can be given by Bozza and Scarpetta (2007)
@@
where $a$ and $b$ are the strong deflection coefficients that are directly associated with the metric functions $A(r_{\mathrm{m}}),B(r_{\mathrm{m}})$ and $C(r_{\mathrm{m}})$. Note that the value of $b$ usually needs to be obtained through numerical integration.
Eq. (33) has been demonstrated to be applicable to photons emitted by sources at any distance outside the event horizon Bozza and Scarpetta (2007). However, its accuracy is compromised when $u$ is considerably smaller than $u_{\mathrm{m}}$, and this is indeed the situation that we will take into consideration later. In comparison, Eq. (27) is only suitable for the case where the source is inside the photon sphere, but gives the higher-order corrections that are omitted by Ref. Bozza and Scarpetta (2007). Therefore, our approximation can be more accurate.
@@ IV.1.2 Comparison with numerical results
From the top panel of Fig. 4 we can observe that when $N=5$, the value of the time delay $\Delta T$ is consistent with $2(\pi-\phi_{\mathrm{S}})u_{\mathrm{m}}$ predicted by Eq. (45). However, $\Delta T$ is not available for $\phi_{\mathrm{S}}=0$. This is because the impact parameter $u|_{n=0}$ is equal to 0 and its corresponding $\epsilon$ defined by (13) becomes divergent, which causes Eq. (37) with $N=5$ to diverge. Such a problem always exists, but $\Delta T$ can be evaluated in the case of $N=0$ due to the relation (41). When $N=5$, the relation (41) is corrected by higher-order terms, and one thus has to calculate $\Delta T$ from Eq. (37) directly.
@@
Furthermore, when calculating the time delay, we have made the assumption that the source, lens and observer lie on a fixed plane, which is just a simple scenario. In more realistic contexts, for instance, when the source maintains a plunging orbit inside the photon sphere Mummery and Balbus (2022), the plane on which the source, lens and observer are located will change over time. How the resulting time delay evolves along the plunging orbit will be our next move.
Finally, black hole mimickers such as ultracompact objects can have an internal structure Cardoso and Pani (2019); Virbhadra and Ellis (2002); Tsukamoto and Harada (2017); Rosa et al. (2022), and photons emitted by the source inside their photon sphere may exhibit different time delay characteristics. How to distinguish black holes and their mimickers through the time delay also deserves future work.